OpenStudy (anonymous):

A test of the prototype of a new automobile shows that the minimum distance for a controlled stop from 94 km/h to zero is 41 m. Find the acceleration, assuming it to be constant as a fraction of the free-fall acceleration. The acceleration due to gravity is 9.81 m/s 2 . Answer in units of g How much time does the car take to stop? Answer in units of s

OpenStudy (anonymous):

convert 94 km/hour to m/s then use \[V _{f}^{2} = V _{o}^{2} +2ad\]

OpenStudy (anonymous):

Have any questions about that?

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

okay 96 converted is 26.11 then v^2=26.11^2+2ad

OpenStudy (anonymous):

yo somebody help